# Diffusion-based Task-oriented Semantic Communications with Model Inversion Attack

### v1

@@ IV-A Module Implementation
The structures of modules involved in our system are as shown in Table I and Table II. We evaluate our proposed system on MNIST and CIFAR10 datasets. Functions of each module are as mentioned above. The compression rates ($\rho$) are set to $8.16\%$ and $6.25\%$ for the MNIST and CIFAR-10 datasets, respectively. These values are chosen merely for the convenience of model training, without any specific or intentional numerical design. Note that before the encoder’s output is sent to the self-denoising module, it undergoes power normalization to make sure that $\sum_{i}(f_{i}^{2}/k)\leq 1$, where $i$ is the pixel index. This operation is performed during the training of the $\mathcal{T}$$-\mathcal{R}$ pair and the training of $\Delta$, to facilitate the denoising process at the receiver. Therefore, the output of the $\mathcal{T}$ does not require further normalization.
@@
Encoder Conv2d+LeakyReLU 16$\times$14$\times$14
Conv2d+LeakyReLU 32$\times$8$\times$8
Conv2d+LayerNorm 1$\times$8$\times$8
Classifier (in Receiver) Linear+ReLU 64$\times$2+1
@@
Linear 1$\times$28$\times$28
TABLE II: Summary of Modules for CIFAR10 Dataset
@@
Encoder Conv2d+LeakyReLU 16$\times$32$\times$32
Conv2d+LeakyReLU 32$\times$16$\times$16
Residual 32$\times$16$\times$16
Conv2d+LeakyReLU +LayerNorm 3$\times$8$\times$8
Classifier (in Receiver) Conv2d+Tanh 64$\times$8$\times$8
Conv2d+Tanh 512$\times$4$\times$4
Residual Block 512$\times$4$\times$4
@@
Strong Classifier (in Third-party) Conv2d+ReLU 64$\times$32$\times$32
Conv2d+ReLU 256$\times$16$\times$16
Residual Block 512$\times$8$\times$8
AdaptiveAvgPool2d 512$\times$2$\times$2
@@
Adversary Conv2d+Tanh 512$\times$8$\times$8
(ConvTrans2d+Tanh +Residual Block)$\times$2 128$\times$32$\times$32
Conv2d+Tanh 3$\times$32$\times$32
@@
Processing Steps Layers Output Dimensions
Encoding ResidualConv Block 128$\times$8$\times$8
Unet Down 128$\times$4$\times$4
Unet Down 256$\times$2$\times$2
Vectorization AvgPool2d+GELU 768$\times$1$\times$1
Embedding 2$\times$×EmbedFC (Time) (256/128)$\times$1$\times$1
2$\times$×EmbedFC (Context) (256/128)$\times$1$\times$1
Decoding ConvTrans2d +GroupNorm+ReLU 256$\times$2$\times$2
UnetUp 128$\times$4$\times$4
UnetUp 128$\times$8$\times$8
Output Conv2d+GroupNorm +ReLU+Conv2d 1$\times$8$\times$8
There are two classifiers in the system, $\Theta$ and stronger $\bar{\Theta}$. When using the MNIST dataset, the input of $\Theta$ is a combination of $\hat{z}$, $\hat{f}$ and $t$, resulting in a dimension of $64\times 2+1$, while the input size of $\bar{\Theta}$ is $1\times 28\times 28$. For the CIFAR10 dataset, the input of $\Theta$ is a concatenation of $\hat{z}$ and $\hat{f}$, with $t$ padded alongside them. Hence, the dimension of the input becomes $(3\times 2)\times 8\times(8+1)=6\times 8\times 9$. The strong classifier $\bar{\Theta}_{\bar{\theta}}$ needs to learn from the combined dataset that includes training dataset $X$, adversary’s dataset $\dot{X}$ and the outputs of the adversary, and aims to achieve as accurate classification as possible. $\bar{\Theta}$ also learns to categorize images that are directly distorted according to channel conditions, denoted as $\ddot{x}$.
The structure of U-Net in the diffusion module is shown in Table III  [^3] Due to limited space and to avoid excessive repetition, here we present the structure of U-Net used in E-DiffSem. For DiffSem, context embedding is not involved, the vectorization step generates 128$\times$2$\times$2, and the output dimensions with 256 should be 128.. We employ U-Net to denoise the input effectively while incorporating the guide information. Specifically, it first processes noisy features by down-sampling them through an encoding step. The down-sampled features are then vectorized and combined with time embeddings and context embeddings (if any) that guide the decoding step. Decoding step up-samples the features and integrates skip connections from the encoding. Finally, convolutional layers reconstruct the denoised features that match the input dimensions. This structure allows the model to effectively denoise images while incorporating time and context information. In our implementation, the diffusion module remains static after training, allowing pre-trained diffusion models to handle and generate transmission information efficiently. This eliminates the need for joint training, significantly lowering both computational overhead and energy usage [37].
Other relevant details are outlined as follows. We use maximum diffusion timestep $T=500$ and $0\leq T^{\prime}\leq T$. We set the learning rate varying from $1\times 10^{-3}$ to $1\times 10^{-4}$. Adam optimizer was utilized for training, along with the ReduceLROnPlateau scheduler to dynamically adjust the learning rate, ensuring efficient convergence and preventing overfitting. Both MNIST and CIFAR10 datasets are divided into two portions, with $60\%$ allocated for the $\mathcal{T}$$-\mathcal{R}$ pair and $40\%$ designated for $\mathcal{A}$.
@@
  • Direct mode (named “Direct”), where images $x\in X$ are directly distorted according to the channel conditions, and then are classified by the strong classifier.
  • Permutation mode (named “Permutation”[40], where the order of elements of images are rearranged according to a specific permutation rule or key, without altering the actual content of the elements. Then permuted images are classified by the strong classifier.
@@ IV-B Results on MNIST dataset
Fig. 4 provides a comprehensive performance comparison of various semantic communication systems for the MNIST dataset with a particular focus on accuracy, under channel SNR 15dB. Here we provide another curve named “Classifier”, which means $\hat{z}$ is directly sent into the classifier $\Theta_{\theta}$ and bypasses the diffusion module. The green solid lines with notations stars and circles represent the classification accuracy of $\hat{z}$ and $\hat{x}$, without other modules involved. They can be simply regarded as a classification task and an image restoration task under different channel conditions that are determined by $t$, and they reveal a fundamental gap that reflects the intrinsic difference between recovering semantic information (task-specific features) and syntactic information (pixel-level details) given specific task and dataset, confirming that task-oriented recovery is inherently more efficient. For example, at $t=300$, the classifier achieves $84\%$ accuracy, while the adversary’s reconstructed images yield only $62\%$ accuracy for the same task. Since semantic information is a highly condensed version of syntax and requires less feature to describe, reconstructing syntax is inherently more demanding than recovering semantics.
@@
Figure 6: Adversary’s output images with different SNR or $t$ under MNIST dataset. 1st row: $-10$ dB; $t=344$. 2nd row: 00 dB; $t=184$. 3rd row: $10$ dB; $t=67$. (a)(e)(i) Original images. (b)(f)(j) IBAL system. (c)(g)(k) IBAL-D system. (d)(h)(l) Our system.
@@
The two red lines at the top of the figure represent the classification accuracy of the receiver after incorporating the diffusion module. It can be observed that using DiffSem provides a slight improvement in the receiver’s recovery, while the label-embedded E-DiffSem significantly enhances system performance. This aligns with the earlier discussion that $\sqrt{\bar{\alpha}_{t}}f_{0}$ can be regarded as an intrinsic guide, and the receiver’s preliminary recovery $\hat{y}$ can serve as guidance to effectively boost diffusion. Most notably, when $t=300$, DiffSem and E-DiffSem improve classification accuracy by 3.35% and 10.03% comparing to the system without diffusion, respectively. One possible explanation for the superior performance of diffusion-based methods lies in their ability to model complex distributions and generate high-quality reconstructions. Diffusion models iteratively refine the output by design, which aligns well with the requirements of semantic communication, where accurate reconstruction of semantic information is crucial. E-DiffSem further optimizes the diffusion process by effectively leveraging the label estimated by the classifier, leading to higher accuracy.
@@
The IBAL and IBAL-D baselines provide an interesting contrast to the proposed methods. Note that the curves in Fig. 4 use SNR as the horizontal axis during training; here they have been recalculated based on Eq. (14) and presented with $t$ as the horizontal axis. We employ the same strong classifier structure for all adversaries including the one in our system, ensuring a consistent ability to assess the amount of task-relevant information in reconstructed images. Since adversarial training, which is grounded in the information bottleneck principle, demands the careful navigation of a dynamic trade-off between the transmission of task-relevant information and the mitigation of adversarial threats, this inevitable compromise inherently limits the amount of information that can be transmitted during standard communication, ultimately causing a pronounced decline in the receiver’s accuracy. In stark contrast, our diffusion-based receiver demonstrates remarkable superiority, achieving significantly higher performance compared to both IBAL systems.
@@
Figure 7: (a)(b)(c) Direct mode and (d)(e)(f) permutation mode, with channel noise $-10$ dB, 00 dB, $10$ dB. Original images are as (a)(e)(i) in Fig. 6.
@@
Fig. 5 shows how accuracy varies with SSIM and PSNR on the MNIST dataset. Here, “Direct” and “Permutation” mean the accuracy of $\bar{\Theta}$’s classification to the original images and permuted images that are directly noised by a certain level of Gaussian white noise  [^4] Here we just use those two baselines as comparisons to show that SSIM or PSNR may be disconnected to the semantic fidelity in some cases, and the appliance of permutation is beyond our scope of this paper. Future work may utilize permutation to further enhance the system performance against model inversion attack. (as shown in Fig. 7), based on Eq. (12); we use a simple permutation method that divides the image into four parts and swaps the position of up-left corner and down-right corner parts. Essentially, our system’s adversary $\mathcal{A}$, as well as the adversaries in the two IBAL systems, can partially achieve the recovery of the input $x$, as evidenced by their SSIM values being higher than those of the direct-noising scenario in most cases with the same accuracy. This aligns with the characteristic of adversaries that have denoising capabilities, as noted in [41]. Compared to the adversary in our system, despite the adversaries in both information-bottleneck-based systems effectively reduce the estimated SSIM and PSNR of the original input obtained by the attacker, we observe that there are cases that the adversary’s classification accuracy can remain relatively high. For instance, the classification accuracy for our system’s attacker (blue dashed line) at SSIM $=0.2$ is $0.47$, while for IBAL-D (purple dashed line with circle markers) at SSIM $=0.175$, it is $0.78$. This indicates that adversarial training merely distorts syntactic features without suppressing semantic information, since the key features remain intact or largely unchanged from the perspective of neural networks. It can be evidenced in Fig. 6 that the adversary’s images in IBAL still have enough semantic information that can be used to categorize, even though their SSIM or PSNR values are lower. The lower plot in Fig. 5, showing the complicated correspondence between PSNR and accuracy, further supports the discussion above, since PSNR ignores perceptual quality and structural information that is important to a classification task [30]. In task-oriented semantic communication, the primary goal is to accurately reconstruct and understand the transmitted task-related information, rather than merely preserving pixel-level fidelity. Therefore, metrics like accuracy, which directly measures the system’s ability to correctly interpret the semantic content, are more appropriate for evaluating the effectiveness of adversarial attacks.
@@
The training strategy of our system prioritizes the recovery of semantic information. The adversary is only tasked with reconstructing images after the $\mathcal{T}-\mathcal{R}$ pair is fully trained to recover task-relevant semantic information, and the amount of information in $\mathcal{T}$’s output can be considered constant. Since we further process the adversary’s output through a strong classifier that effectively converting the image reconstruction task into a classification task, it allows us to evaluate how much of the “task-relevant” information is contained within the reconstructed images, which is directly tied to the quality of the image reconstruction. However, due to the limited information provided by $\mathcal{T}$ and the fact that the $\mathcal{T}$’s output is not primarily designed for image reconstruction, the adversary’s ability to recover syntactic information is inherently biased, resulting in less “task-relevant” information embedded in the reconstructed images. If the $\mathcal{T}-\mathcal{A}$ pair is trained first, followed by the receiver $\mathcal{R}$, the training process would inevitably prioritize a “high-information task” (image reconstruction) before transitioning into a “low-information task” (classification). This approach would likely lead to the encoder producing excessive information, which, according to the evaluation standard defined in Def. (1), would benefit the adversary by providing more access to task-relevant semantic information, even if the recovered images have relatively low SSIM (or PSNR) values. Consequently, this strategy would compromise the privacy protection intended for the classifier.
@@
Fig. 8 illustrates the robustness of our DiffSem system across diverse channel conditions. As outlined in Sec. III Part C, the transmitter dynamically adjusts the level of self-noising pre-added noise based on the estimated channel condition. When facing high SNR regimes, the self-noising module becomes the primary driver of signal distortion, meaning that variations in $t$ more directly influence the performance of both the receiver and the attacker, allowing precise control via $t$. Conversely, in poor channel conditions, the wireless channel itself dominates, allowing the self-noising module to flexibly control the quality of the transmitted signal while preventing over-compression to exert minimal impact on the signal state at the receiver. For instance, at a channel SNR of $-5$ dB, the dominance of channel noise results in a nearly constant information gap between the classifier and the adversary. We observe that for $t>250$, the performance of both the receiver and the adversary slightly deteriorates compared to scenarios with higher channel SNR. This discrepancy may stem from the discrete nature of $\bar{\alpha}_{t}$, causing deviations between practical estimates and theoretical calculations. Overall, our self-noising module effectively modulates the degree of signal distortion in response to wireless channel conditions, maintaining an appropriate noise margin and ensuring reliable operation under both high and low SNR environments.
IV-C Results on CIFAR10 dataset
Figure 9: System performances using CIFAR10 dataset.

### v2

@@ IV-A Module Implementation
The structures of modules involved in our system are as shown in Table I and Table II. Functions of each module are as mentioned above. The compression rates ($\rho$) are set to $8.16\%$ and $6.25\%$ for the MNIST and CIFAR-10 datasets, respectively. These values are chosen merely for the convenience of model training, without any specific or intentional numerical design. Note that before the encoder’s output is sent to the self-noising module, it undergoes power normalization to make sure that $\sum_{i}(f_{i}^{2}/k)\leq 1$, where $i$ is the pixel index. This operation is performed during the training of the $\mathcal{T}$-$\mathcal{R}$ pair and the training of $\Delta_{\delta}$, to facilitate the denoising process at the receiver. Therefore, the output of the $\mathcal{T}$ does not require further normalization.
@@
[…]
@@
[…]
TABLE II: Summary of Modules for CIFAR-10 Dataset
@@
[…]
@@
[…]
@@
[…]
@@
Steps Layers Output Dimensions
[…]
There are two classifiers in the system, $\Theta_{\theta}$ and stronger $\bar{\Theta}_{\bar{\theta}}$. When using the MNIST dataset, the input of $\Theta_{\theta}$ is a combination of $\hat{z}$, $\hat{f}$ and $t$, resulting in a dimension of $64\times 2+1$, while the input size of $\bar{\Theta}_{\bar{\theta}}$ is $1\times 28\times 28$. For the CIFAR-10 dataset, the input of $\Theta_{\theta}$ is a concatenation of $\hat{z}$ and $\hat{f}$, with $t$ padded alongside them. Hence, the dimension of the input becomes $(3\times 2)\times 8\times(8+1)=6\times 8\times 9$. The strong classifier $\bar{\Theta}_{\bar{\theta}}$ needs to learn from the combined dataset that includes $x$, $\hat{x}$ and $\dot{x}$, and aims to achieve as accurate classification as possible. $\bar{\Theta}_{\bar{\theta}}$ also learns to categorize images that are directly distorted according to channel conditions, denoted as $\ddot{x}$.
[…]
Other relevant details are outlined as follows. We use maximum diffusion timestep $T=500$ and $0\leq T^{\prime}\leq T$. We set the learning rate varying from $1\times 10^{-3}$ to $1\times 10^{-4}$. Adam optimizer was utilized for training, along with the ReduceLROnPlateau scheduler to dynamically adjust the learning rate, ensuring efficient convergence and preventing overfitting. Both MNIST and CIFAR-10 datasets are divided into two portions, with $60\%$ allocated for the $\mathcal{T}$-$\mathcal{R}$ pair and $40\%$ designated for $\mathcal{A}_{a}$.
@@
  • Direct mode (named “Direct”), where images $x\in X$ are directly distorted according to the channel conditions to get $\ddot{x}$, and then are classified by the strong classifier.
  • Permutation mode (named “Permutation”) [31], where the order of elements of images are rearranged according to a specific permutation rule or key, without altering the actual content of the elements. Then permuted images are classified by the strong classifier.
@@ IV-B Results on MNIST dataset
Fig. 4 provides a comprehensive performance comparison of various semantic communication systems for the MNIST dataset with a particular focus on accuracy, under channel SNR 15 dB. Here we provide another curve named “Classifier”, which means $\hat{z}$ is directly sent into the classifier $\Theta_{\theta}$ and bypasses the diffusion module. The green solid lines with notations stars and circles represent the classification accuracy of $\hat{z}$ and $\hat{x}$, without other modules involved. They can be simply regarded as a classification task and an image restoration task under different channel conditions that are determined by $t$, and they reveal a fundamental gap that reflects the intrinsic difference between recovering task-specific features (task-oriented semantic information) and pixel-level details (syntactic information) given the specific task and dataset [8], confirming that task-oriented recovery is inherently more efficient. For example, at $t=300$, the classifier achieves $84\%$ accuracy, while the adversary’s reconstructed images yield only $62\%$ accuracy for the same task. Since semantic information is a highly condensed version of syntax and requires less feature to describe, reconstructing syntax is inherently more demanding than recovering semantics.
@@
Figure 6: Adversary’s output images with different SNR or $t$ under MNIST dataset. 1st row: 10-10 dB; $t=344$. 2nd row: 0 dB; $t=184$. 3rd row: $10$ dB; $t=67$. (a)(b)(c) Original images. (d)(e)(f) IBAL system. (g)(h)(i) IBAL-D system. (j)(k)(l) Our system.
@@
The two red lines at the top of the figure represent the classification accuracy of the receiver after incorporating the diffusion module. It can be observed that using DiffSem provides a slight improvement in the receiver’s recovery, while the label-embedded E-DiffSem significantly enhances system performance. This aligns with the earlier discussion that $\sqrt{\bar{\alpha}_{t}}f_{0}$ can be regarded as an intrinsic guide, and the receiver’s preliminary recovery $\hat{y}$ can serve as guidance to effectively boost diffusion. Most notably, when $t=300$, DiffSem and E-DiffSem improve classification accuracy by 3.35% and 10.03% compared to the system without diffusion, respectively. One possible explanation for the superior performance of diffusion-aided methods lies in their ability to model complex distributions and generate high-quality reconstructions. Diffusion models iteratively refine the output by design, which aligns well with the requirements of semantic communication, where accurate reconstruction of semantic information is crucial. E-DiffSem further optimizes the diffusion process by effectively leveraging the label estimated by the classifier, leading to higher accuracy.
@@
The IBAL and IBAL-D baselines provide an interesting contrast to the proposed methods. Note that the curves in Fig. 4 use SNR as the horizontal axis during training; here they have been recalculated based on Eq. (13) and presented with $t$ as the horizontal axis. We employ the same strong classifier structure for all adversaries including the one in our system, ensuring a consistent ability to assess the amount of task-relevant information in reconstructed images. Since adversarial training, which is grounded in the information bottleneck principle, demands the careful navigation of a dynamic trade-off between the transmission of task-relevant information and the mitigation of adversarial threats, this inevitable compromise inherently limits the amount of information that can be transmitted during standard communication, ultimately causing a pronounced decline in the receiver’s accuracy. In stark contrast, our diffusion-aided receiver demonstrates remarkable superiority, achieving significantly higher performance compared to both IBAL systems.
@@
Figure 7: (a)(b)(c) Direct mode and (d)(e)(f) permutation mode, with channel noise 10-10 dB, 0 dB, $10$ dB. Original images are as (a)(b)(c) in Fig. 6.
@@
Fig. 5 shows how accuracy varies with SSIM and PSNR on the MNIST dataset. Here, “Direct” and “Permutation” mean the accuracy of $\bar{\Theta}_{\bar{\theta}}$’s classification to the original images and permuted images that are directly noised by a certain level of Gaussian white noise  [^6] Here we just use those two baselines as comparisons to show that SSIM or PSNR may be disconnected to the semantic fidelity in some cases, and the application of permutation is beyond our scope of this paper. Future work may utilize permutation to further enhance the system performance against model inversion attack. (as shown in Fig. 7), based on Eq. (11); we use a simple permutation method that divides the image into four parts and swaps the position of up-left corner and down-right corner parts. Essentially, our system’s adversary $\mathcal{A}_{a}$, as well as the adversaries in the two IBAL systems, can partially achieve the recovery of the input $x$, as evidenced by their SSIM values being higher than those of the direct-noising scenario in most cases with the same accuracy. This observation is consistent with neural-based systems that possess denoising capabilities [47]. Compared to the adversary in our system, despite the adversaries in both information-bottleneck-based systems effectively reduce the estimated SSIM and PSNR of the original input obtained by the attacker, we observe that there are cases that the adversary’s classification accuracy can remain relatively high. For example, when the SSIM is 0.2, the attacker in our system (blue dashed line) achieves a accuracy of 0.47, whereas IBAL-D (purple dashed line with circle markers) reaches 0.78 at SSIM =0.175=0.175. This indicates that adversarial training merely distorts syntactic features without suppressing semantic information, since the key features remain intact or largely unchanged from the perspective of neural networks. It can be evidenced in Fig. 6 that the adversary’s images in IBAL still have enough semantic information that can be used to categorize, even though their SSIM or PSNR values are lower. The lower plot in Fig. 5, showing the complicated correspondence between PSNR and accuracy, further supports the discussion above, since PSNR ignores perceptual quality and structural information that is important to a classification task [32]. In task-oriented semantic communication, the primary goal is to accurately reconstruct and understand the transmitted task-related information, rather than merely preserving pixel-level fidelity. Therefore, metrics like accuracy, which directly measures the system’s ability to correctly interpret the semantic content, are more appropriate for evaluating the effectiveness of adversarial attacks.
@@
The training strategy of our system prioritizes the recovery of semantic information. The adversary is only tasked with reconstructing images after the $\mathcal{T}$-$\mathcal{R}$ pair is fully trained to recover task-relevant semantic information, and the amount of information in $\mathcal{T}$’s output can be considered constant. Since we further process the adversary’s output through a strong classifier that effectively converting the image reconstruction task into a classification task, it allows us to evaluate how much of the “task-relevant” information is contained within the reconstructed images, which is directly tied to the quality of the image reconstruction. However, due to the limited information provided by $\mathcal{T}$ and the fact that the $\mathcal{T}$’s output is not primarily designed for image reconstruction, the adversary’s ability to recover syntactic information is inherently biased, resulting in less “task-relevant” information embedded in the reconstructed images. If the $\mathcal{T}-\mathcal{A}$ pair is trained first, followed by the receiver $\mathcal{R}$, the training process would inevitably prioritize a “high-information task” (image reconstruction) before transitioning into a “low-information task” (classification). This approach would likely lead to the encoder producing excessive information, which would benefit the adversary by providing more access to task-relevant semantic information, even if the recovered images have relatively low SSIM (or PSNR) values. Consequently, this strategy would compromise the privacy protection intended for the classifier.
@@
Fig. 8 illustrates the robustness of our DiffSem system across diverse channel conditions. As outlined in Sec. III Part C, the transmitter dynamically adjusts the level of self-noising pre-added noise based on the estimated channel condition. When facing high SNR regimes, the self-noising module becomes the primary driver of signal distortion, meaning that variations in $t$ more directly influence the performance of both the receiver and the attacker, allowing precise control via $t$. Conversely, in poor channel conditions, the wireless channel itself dominates, allowing the self-noising module to flexibly control the quality of the transmitted signal while preventing over-compression to exert minimal impact on the signal state at the receiver. For instance, at a channel SNR of 5-5 dB, the dominance of channel noise results in a nearly constant information gap between the classifier and the adversary. We observe that for $t>250$, the performance of both the receiver and the adversary slightly deteriorates compared to scenarios with higher channel SNR. This discrepancy may stem from the discrete nature of $\bar{\alpha}_{t}$, causing deviations between practical estimates and theoretical calculations. Overall, our self-noising module effectively modulates the degree of signal distortion in response to wireless channel conditions, maintaining an appropriate noise margin and ensuring reliable operation under both high and low SNR environments.
IV-C Results on CIFAR-10 dataset
Figure 9: System performances using CIFAR-10 dataset.